OpenStudy (anonymous):

Relative to the ground below, how many joules of PE does a 700 N boulder have at the top of a 2 m ledge? PE=1400 J KE=1400 J What will be its speed on impact?

OpenStudy (mrnood):

Kinetic energy is calculated as KE=1/2 MV^2 (half m times v squared) You need to know the MASS of the boulder. The first part of the question gives you the weight (700N). If we assume (reasonably) that the boulder is on earth then the weight = m*g g is approximately 9.81m/s^2 - so you can calculate the mass. So you know the KE =1400J KE = 1/2 mv^2 so you can calculate v

OpenStudy (wolfe8):

PE=mgh where mg is the weight that you are given. So you basically multiply the weight with the height. This PE will be transformed into KE, meaning your KE=PE. And KE=\[\frac{ 1 }{ 2 }mv ^{2}\] You can get m by using the formula w=mg. Then you will just have to fine the velocity.